# Data-Based System Representation and Synchronization for Multiagent Systems

### v1

@@ IV DATA-BASED SYNCHRONIZATION OF HETEROGENEOUS SYSTEMS
In this section, we propose to use a dynamic controller as in (39)-(40) designed entirely from measured data. A discrete-time version of this problem was studied in [12], where the equations (41)-(42) are solved by leveraging data matrices. However, that solution requires partial knowledge of the systems model. In particular, the matrices $C_{i}$ must be known for all agents $i$, and complete knowledge of the leader dynamics (16) is needed. In this paper, we provide a solution that does not require any knowledge of the systems models (15)-(16).
@@ IV-A Data-based representation of heterogeneous dynamics
First, we consider $N$ dynamical systems with (potentially) different linear models as in (15). Since the models are different, data must be collected from each agent individually. Thus, assuming that all pairs $(A_{i},B_{i})$ are controllable, apply a PCPE input of order $n_{i}+1$ to each agent $i$, and collect the data $(u_{i},x_{i},y_{i})$ (see Remark 2 regarding the availability of both state and output data). Now, using (5) and (8), we represent every input-state-output trajectory $(\bar{u}_{i},\bar{x}_{i},\bar{y}_{i})$ of agent $i$ as
$\begin{bmatrix}\mathcal{H}_{T}(u_{i})\\ \mathcal{H}_{T}(x_{i}(t))\\ \mathcal{H}_{T}(y_{i}(t))\end{bmatrix}\alpha_{i}(t)=\begin{bmatrix}\bar{u}_{i}(t)\\ \bar{x}_{i}(t)\\ \bar{y}_{i}(t)\end{bmatrix},$ (43)
for $0\leq t<T$.
For the leader (16), we must assume that we measure a state trajectory $x_{0}$ such that $\operatorname*{rank}(\mathcal{H}_{T}(x_{0}(t)))~{}=~{}n_{0}$ for all $0\leq t<T$, without the need of a persistently exciting input [^2] Compare this assumption with the results in [24], where the conditions of the discrete-time Willems’ lemma for data-based representation of system trajectories were relaxed for the case of uncontrollable systems.. Using this state information, together with the corresponding output data from (16b), every leader state-output trajectory is expressed as
@@ V SIMULATIONS
In this section, we apply the results of Theorem 3 to achieve output synchronization in a heterogeneous multiagent system (15)-(16). In particular, we consider one leader and four follower agents connected in the communication graph displayed in Figure 1. If agent $i$ receives information from agent $j$, then we set $a_{ij}=1$. Since only agent $1$ receives information from the leader, we have $g_{1}=1$, while $g_{i}=0$ for $i=2,3,4$.
001111222233334444
Figure 1: Communication graph for simulation.
Here we consider the same system models for the agents as those used in [3, Section 4]. That is, for the leader (16) we use the matrices
@@ VI CONCLUSIONS
This paper presented novel data-based solutions to the synchronization problem in multiagent systems. First, the problem with identical agents was considered and a solution using linear matrix inequalities was proposed. This procedure yields distributed static controllers for each agent, where no dynamic variables are required to obtain stability. Then, the output synchronization problem was solved for heterogeneous systems. This was achieved by means of dynamic control inputs that were obtained using measured data and without any knowledge of the leader or the followers dynamics.
Future lines of research include investigating conditions to achieve synchronization in heterogeneous agents using LMIs as in (38), as well as developing distributed controllers that are robust against noise in the data measurements.

### v2

@@ IV DATA-BASED SYNCHRONIZATION OF HETEROGENEOUS SYSTEMS
In this section, we propose to use a dynamic controller as in (38)-(39) designed entirely from measured data. A discrete-time version of this problem was studied in [12] and [13], where the equations (40)-(41) are solved by leveraging data matrices. However, those solutions require partial knowledge of the systems model. In particular, the matrices $C_{i}$ must be known for all agents $i$, and complete knowledge of the leader dynamics (16) is needed. In this paper, we provide a solution that does not require any knowledge of the systems models (15)-(16).
@@ IV-A Data-based representation of heterogeneous dynamics
First, we consider $N$ dynamical systems with (potentially) different linear models as in (15). Since the models are different, data must be collected from each agent individually. Thus, assuming that all pairs $(A_{i},B_{i})$ are controllable, apply a PCPE input of order $n_{i}+1$ to each agent $i$, and collect the data $(u_{i},x_{i},y_{i})$ (see Remark 2 regarding the availability of both state and output data). Now, using (5) and (8), we represent every input-state-output trajectory $(\bar{u}_{i},\bar{x}_{i},\bar{y}_{i})$ of agent $i$ for $0\leq t<T$ as
$\begin{bmatrix}\mathcal{H}_{T}(u_{i})\\ \mathcal{H}_{T}(x_{i}(t))\\ \mathcal{H}_{T}(y_{i}(t))\end{bmatrix}\alpha_{i}(t)=\begin{bmatrix}\bar{u}_{i}(t)\\ \bar{x}_{i}(t)\\ \bar{y}_{i}(t)\end{bmatrix}.$ (42)
For the leader (16), we must assume that we measure a state trajectory $x_{0}$ such that $\operatorname*{rank}(\mathcal{H}_{T}(x_{0}(t)))~{}=~{}n_{0}$ for all $0\leq t<T$, without the need of a PE input [^2] Compare this assumption with the results in [25], where the conditions of the discrete-time Willems’ lemma for data-based representation of system trajectories were relaxed for the case of uncontrollable systems.. Using this state information and the corresponding output data from (16b), every leader state-output trajectory is expressed as
@@ V SIMULATIONS
In this section, we apply the results of Theorem 3 to achieve synchronization in a heterogeneous multiagent system [^3] Matlab code available https://doi.org/10.25835/wj72f6ba. (15)-(16). We consider one leader and four follower agents connected in the communication graph displayed in Figure 1. If agent $i$ receives information from agent $j$, then we set $a_{ij}=1$. Since only agent $1$ receives information from the leader, we have $g_{1}=1$, while $g_{i}=0$ for $i=2,3,4$. Here we consider the same system models as those used in [3, Section 4]. That is, for the leader (16) we use the matrices
@@ VI CONCLUSIONS
This paper presented novel data-based solutions to the synchronization problem in multiagent systems. First, the problem with identical agents was considered and a solution using LMIs was proposed. This procedure yields distributed static controllers for each agent, where no dynamic variables are required. Then, the output synchronization problem was solved for heterogeneous systems by means of dynamic control inputs without any knowledge of the leader or the followers dynamics. As future work we will investigate conditions to achieve synchronization in heterogeneous agents using LMIs as in (37), and develop distributed controllers that are robust against noise in the data measurements.